1. The machine has been wonderful this morning. We got 6.5 of great data. 2. I took a number of production runs, and I took two zero field runs. 3. I noticed that Tof 2 still has problems. I think this has been noticed before, but I'm not sure what is to be done about it. Also, the default values for a number of HV channels (see shift check lists) do not agree with the target values. Also, the tile signals that KH noticed were missing are missing because these channel are not bias. I'm assuming this is because the detector element is inactive. _______________________________________________ Brahms-l mailing list Brahms-l@lists.bnl.gov http://lists.bnl.gov/mailman/listinfo/brahms-lReceived on Wed Jan 26 07:42:58 2005
This archive was generated by hypermail 2.1.8 : Wed Jan 26 2005 - 07:43:19 EST